### v1

# Network of velocity-coherent filaments formed by supersonic turbulence in a very-high-velocity HI cloud

Network of velocity-coherent filaments formed by supersonic turbulence in a very-high-velocity HI cloud
The warm neutral medium (WNM) was thought to be subsonically/transonically turbulent, and it lacks a network of intertwined filaments 1, 2 that are commonly seen in both molecular clouds (refs. 3, 4, 5, 6, 7, 8, 9, 10) and cold neutral medium (CNM) (refs.11, 12, 13). Here, we report Hi 21 cm line observations of a very-high-velocity (-330 km s-1 -250 km s-1) cloud (VHVC), using the Five-hundred-meter Aperture Spherical radio Telescope (FAST) with unprecedented resolution and sensitivity. For the first time, such a VHVC is clearly revealed to be a supersonic WNM system consisting of a network of velocity-coherent Hi filaments. The filaments are in the forms of slim curves, hubs, and webs, distributed in different layers within the position-position-velocity (ppv) data cube. The entire cloud has skewed log-normal probability distribution of column density and the filaments themselves show asymmetrical radial density profiles, indicating shock compression by supersonic magnetohydrodynamic (MHD) turbulence, as is also confirmed by our MHD simulation (sonic Mach number and Alfvén Mach number ). This work suggests that hierarchical filaments can be established by shocks in a low-density WNM, where gravity is negligible, offering a viable pathway to structure formation in the earliest evolutionary phases of the interstellar medium (ISM).
Main body
Hierarchical density structures in the warm neutral medium (WNM) have not yet been well characterized by observations even though some theoretical works have predicted that elongated low-density parallel filaments can be formed in the sub-Alfvénic WNM 14. High-velocity clouds (HVCs; km s-1) that contain a significant amount of WNM gas and locate in relatively isolated environments provide an excellent laboratory to explore this issue. HVCs are Hi clouds at radial velocities incompatible with the regular rotation of the Galactic disc15, 16, and the absorption lines of distant halo stars indicate that they are generally located several kpc above the Galactic plane (refs.17, 18, 19, 20). The overall negative velocities of HVCs (refs.16, 18, 21) suggest that they may trace the infall of early-phase low-metalicity material onto the plane of the Milky Way (refs.22, 23, 24, 25). HVCs under the view of the all-sky Hi surveys such as HI4PI21 are in the form of atomic gas streams or complexes, because of limited resolution (; see also Extended Data Fig. 1). A few HVCs with km s-1 that were observed by the GALFA-Hi survey of Arecibo 26 have been revealed to be WNM systems consisting of slim filaments27, 28. Most of those filaments are aligned in parallel patterns, and may be regulated by the global dynamics of the gas streams, or by the dynamic processes in the Galactic plane 28, 29. In contrast, the very-high-velocity clouds (VHVCs; km s-1) are for certain not coupled with the Galactic disc, and may have a different origin from common HVCs (refs. 29, 30, 31, 32). Detailed studies of VHVCs, however, are extremely rare in the past. Observations of the Hi 21 cm line (, =1-0; MHz)15 with sufficiently high resolutions to reveal the internal morphology and gas dynamics of VHVCs are essential for studying their origins and the underlying physical processes involved in the formation of hierarchical gas structures within the WNM.
The very-high-velocity cloud G165
The HI4PI survey (refs.33, 34, 35, 36) provided so far the most detailed Hi images of all-sky HVCs21. VHVCs are dominated by the Magellanic Stream (MS) 37, 38 and the anticentre (AC) complex (refs.39, 40, 29). The VHVC component of the AC complex, denoted as G165 VHVC in this work, is the most notable one in brightness and size on the sky (Extended Data Fig. 1), which stands out as an ideal object for investigating filament formation in extreme environments. Besides the VHVC component (-330 km s-1 -250 km s-1), at the line of sight of G165 there is also an HVC component (-140 km s-1 -80 km s-1) and a foreground component (-20 km s-1 20 km s-1). G165 HVC is part of the Cohen Stream, which was measured to have a distance bracket of 5.0 to 11.7 kpc18. In contrast, the H brightness of G165 VHVC implies a distance () on the order of 15 kpc18, the value we adopted. The G165 VHVC is gravitationally unbound with a size of 100 pc, and a total gas mass of (see methods). Limited by poor angular resolution (), the inner structures of the G165 VHVC were not resolved but did show clumpy features in the HI4PI survey.
A WNM system with supersonic dynamics
We mapped the G165 VHVC using the Five-hundred-meter Aperture Spherical radio Telescope (FAST), with unprecedented angular resolution (), velocity resolution ( km s-1), and line sensitivity (60 mK per channel). G165 VHVC is sufficiently resolved in the integrated intensity map of the Hi 21 cm line (left panel in Fig. 1). Different velocity subcomponents of the G165 VHVC are spectrally blended, but can be well separated by careful Gaussian decomposition (Fig. 1; see Sect. Methods for details). These subcomponents are distributed in several layers in the position-position-velocity (ppv) cube (Fig. 2). The layers are interlaced with each other, and show large-scale velocity fluctuations (or velocity wiggles) as also witnessed in molecular interstellar medium 41. The central velocities () of the subcomponents are divided into five groups ranging from km s-1 to km s-1 (Fig. 3a), with no one-to-one correspondence between the velocity groups and gas layers.
The line widths (full width at half maximum; FWHM) of the velocity subcomponents are slightly larger than 20 km s-1 (Fig. 3c), the value of the typical thermal line widths of the WNM42, 43, 2. Assuming a gas kinetic temperature of the G165 VHVC of 6000 K, the typical value of WNM2, the thermal dispersion is estimated as 7 km s-1, yielding the transonic () nature of individual velocity subcomponents. However, different velocity groups are separated by 20 km s-1, suggesting the whole cloud could be in supersonic turbulence. Considering that different gas layers are intersected, forming a dynamically correlated system (see also the next sections), all velocity groups can be treated as a whole system. The total dispersion of the centroid velocities (), a good tracer of the internal turbulent motions within a cloud, is up to 20 km s-1, indicating a global of 3. We note that the observed distribution of line widths (Fig. 3c) can be well reproduced by our MHD simulation of supersonic and trans-Alfvénic gas ( and ). The same simulation also spans a large velocity range with several velocity groups (Fig. 3b), as the same in observations.
The median values of column density ( cm-2) and volume density (0.2 cm-3) of the G165 VHVC (Sect. Methods) are also similar to the typical values of the WNM42. The column density probability distribution function (-PDF) of the decomposed subcomponents is higher-end suppressed (or lower-end enhanced), and deviates from a lognormal shape (Fig. 3d). The lognormal behavior of an -PDF is usually explained by supersonic turbulence44, 45, 46, but the intermittency of a supersonic system caused by head-on collisions of strongly shocked gas could lead to considerable deviations47, 48. The observed -PDF can also be depicted by our MHD simulation models of supersonic and trans-Alfvénic gas ( and ; see below).
To conclude, we consider the G165 VHVC as a WNM system with supersonic dynamics. This is significantly different from the majority of the WNM in Galactic plane or in HVCs (e.g., Smith cloud), where turbulence is subsonic to transonic 14, 28, 49, 46.
Web-like filamentary structures
Filamentary structures in the G165 region, which cannot be resolved by HI4PI, are clearly seen by FAST on the Hi moment 0 maps of its VHVC, HVC, and foreground components (Extended Data Fig. 2). The HVC components show parallel filaments elongated along the Cohen Stream, similar to the famous Smith HVC 28. In contrast, both the foreground and VHVC components show web-like filamentary structures. The foreground component has no clear boundary, while the VHVC shows clear boundaries defined by long and slim filaments. Another difference between the foreground and VHVC components is that the filaments tend to be in parallel mode on the channel maps of the foreground component, but stay intersected at different velocity slices of the VHVC components (Extended Data Fig. 2). These comparisons indicate that the filamentary system of the G165 VHVC is really web-like rather than in parallel mode, which is substantially different from those in typical HVCs (e.g., Smith HVC 28) and low/intermediate velocity Hi clouds 11, 13.
Filamentary structures in the G165 VHVC are visible in different layers of the velocity subcomponents (Fig. 2). However, this may be misleading due to projection effects, which enhances the surface intensity at places tangent to the line of sight. We developed a tool, ClimbRidge, based on a density ridge algorithm50, 51, to extract the density ridges of G165 VHVC in both 2-d and 3-d cases (Sect. Methods). ClimbRidge gives similar results as the well known tool FilFinder in 2-d case (Extended Data Fig. 3), and can be naturally applied to 3-d (ppv) data to trace truly velocity-coherent filaments (Fig. 2 and Extended Data Fig. 4). Minimal Spanning Trees (MSTs) of the points on the density ridges were built (Extended Data Fig. 6; see Sect. Methods), and the lengths of filament segments were measured (Fig. 3e). As one can see, the distribution of segment lengths in our MHD simulation of supersonic and trans-Alfvénic gas ( and ) remarkably resembles the observed one.
The filaments in the G165 VHVC are mainly distributed in three layers in ppv space (denoted as L1, L2 and L3 in Extended Data Fig. 4). Each layer has an angular size larger than 2.5∘. On the plane of (RA, DEC), the G165 VHVC can be divided into its eastern, western, and southern parts. L1 is the main body of the western part. L2 spans from the eastern part to the western part, with a global velocity gradient of 0.2 km s-1 per arcmin (0.05 km s-1 pc-1). L3 extends from the western part to the southern part, with a global velocity gradient of 0.3 km s-1 per arcmin (0.07 km s-1 pc-1) (Extended Data Fig. 4). The gas flares of L2, with their velocities spreading to km s-1, also has a filamentary structure. The filaments of the G165 VHVC are intersected with each other, building a filamentary network in the ppv space.
There are long and slim filaments that go through different layers, indicating that different layers are indeed not independent and form a dynamically correlated system. Three representative filaments (denoted as B1, B2, and B3) are highlighted in the right panel of Fig. 2 and Extended Data Fig. 4. We view these long filaments from different viewing angle in the ppv space (Extended Data Fig. 5), confirming that the density ridges do trace the skeletons of the filamentary structures in the ppv space. Basic parameters of the three filaments are fitted and listed in Extended Data Table 1. The velocity gradient of the filaments is comparable with that of the gas layers (0.1 km s-1 pc-1), contributing a velocity span of km s-1. This also suggests that the filaments in G165 VHVC are supersonic. The filaments of the G165 VHVC tend to be radially asymmetrical on the - plane (for example see Extended Data Fig. 5). The intensity difference between the two sides of the filament is comparable with the intensity enhancement of the central ridge (Extended Data Fig. 7). Our MHD simulations suggest that the asymmetrical density enhancement of filaments could be induced by supersonic turbulence (Fig. 3f).
Overall, we suggest that: (1) The G165 VHVC is a supersonic WNM system; (2) A filamentary network has already been established in such an early-phase atomic system; (3) The dynamic structure (several gas layers; multiple velocity groups; skewed -PDF) and the filamentary network (including their asymmetrical radial profiles) could be induced by supersonic shocks. The effect of self-gravity of the G165 VHVC onto itself and its filamentary substructures is negligible (Sect. Methods). To conclude, filaments, and even web-like filamentary structures, can be well established in very low density but supersonic turbulent gas clouds (e.g., WNM in G165 VHVC), where self-gravity is not important.
Physical environment implication through numerical simulation
To further check the turbulent levels in the G165 VHVC, we conducted a series of isothermal MHD simulations with subsonic (), transonic () and supersonic () turbulence (Sect. Methods). We do not have magnetic field information of the region but we expect that magnetic fields will have effects on the partially ionized WNM. Therefore, we also perform a parameter study with 0.5, 1, and 2 as explained in the Methods section. Slim filaments with asymmetrical profiles can be clearly produced by supersonic turbulence combined with trans/super-Alfvénic numbers (Extended Data Fig. 7). In ppv space, the simulated data shows a multi-layered pattern as in the observations (Supplementary Fig. 4). The velocity gradient and velocity wiggles seen in the observed gas layers (Fig. 2) can also be reproduced (Supplementary Fig. 4). We argue that the intermittency of supersonic turbulence could be responsible for the asymmetrical profiles of filaments (Fig. 3f), which further contribute to the lower-end enhancement of the PDF (Fig. 3d).
Gaussian decomposition and filament extracting procedures were applied to the simulated ppv cubes as to the observed one (Sect. Methods). The three subsonic and transonic models () have a much narrower -PDF than the supersonic models and the observed -PDF (Extended Data Fig. 8). This indicates that we need a supersonically turbulent region to recreate the -PDF of the observed region. The two slightly super-Alfvénic models () have longer low density tails than the observed -PDF. The best fit to the observation is the model with and . The best-fit simulation can also statistically reproduce the observed distribution of the line width, central velocity, and column density of the decomposed subcomponents (Figure 3a–d). It again supports that the G165 VHVC is a dynamically correlated supersonic and trans-Alfvénic system. We extracted filaments in the ppv space of the best-fit model and successfully reproduced the observed distribution of the lengths of the filament segments (Figure 3e). The filament network in G165 VHVC is most likely formed due to supersonic MHD turbulence.
The magnetic field strength of the G165 VHVC is then estimated to be (see Sect. Methods). However, the lack of the low-density tail of the observed -PDF could also be a result of detection sensitivity. Therefore, our simulations do not completely rule out the cases of somewhat weaker magnetic fields. We do not know how the G165 VHVC is bounded (or not). Besides a magnetic field, ram pressure and bulk motion may also contribute to anisotropic dynamics that have not been considered by our simulation. At present, we can not exclude effects due to these anisotropic dynamics. Instead, our comparisons suggest that the supersonic MHD turbulence, in the sense of a zero-order approach, can explain the multi-layer pattern and filament-network formation of G165 VHVC.
Analog to filaments in Molecular Clouds
Hierarchical filamentary structures are ubiquitous in Galactic molecular clouds (MCs), in the form of slim curves52, hubs53, and webs3. Velocity gradients along the filaments are typically km s-1 pc-1 in MCs54, 7, and can reach 5 km s-1 pc-1 within massive molecular clumps8. It was often interpreted as gas flow along the filament under the effect of gravity. This work implies that velocity gradients along the filaments have been established in early-phase atomic clouds like the G165 VHVC, and will increase from 0.1 km s-1 pc-1 to several km s-1 pc-1 as their natal clouds vary from hundred-pc-scale atomic clouds to sub-pc-scale molecular clumps. The contribution of turbulence to the velocity gradients along the filaments should not be ignored. In spite of the possible different origins of velocity gradients, the filaments of the G165 VHVC could also play a role in transporting material between different layers (Extended Data Fig. 4). The web-like filaments within G165 VHVC do not just intersect with each other. Instead, different filaments can intersect each other constituting hub-like structures with mass assembling at the intersection knots. For example, the long filament B1 and its short fibers form such a hub-filament system (Extended Data Fig. 5). To exclude projection bias, we changed the viewing angles to find out the fibers that are truly connected to B1. Three such subbranches are found, each of which contributes an intensity enhanced hub at the intersection point. In a word, all filamentary types (slim curves, hubs, and webs) that commonly exist in MCs are also revealed in the G165 VHVC, indicating that the origins of the filaments in these two distinctly different types of ISMs may be regulated by some shared physical mechanisms including, most likely, supersonic MHD turbulence. Observations of more VHVCs in the future should yield a deeper understanding of the physical nature of VHVCs and their formation.
Methods
1. FAST observation and data reduction
The observations were conducted in on-the-fly (OTF) mode using the Five-hundred-meter Aperture Spherical radio Telescope (FAST)55. The illuminated aperture () of FAST is 300 meter for zenith angle 26.4∘, and decreases by one third the value for 56. The tracking accuracy is 0.2′. The -band (1–1.5 GHz) front end is a 19-beam dual-polarization (XX and YY) receiver (Supplementary Fig. 1). It has an FWHM beam size () of 3′ at 1.4 GHz, a system temperature of less than 24 K 56, and a beam efficiency for extended sources of 0.85 56, 57.
The spectral backend provides 1024k (about one million) channels, providing a velocity resolution of 0.1 km s-1 at 1.4 GHz, much better than that of the HI4PI survey ( km s-1)36. The analog-to-digital converter (ADC) of the FAST samples the data every 4 ns. The time intervals related to the spectral backend are thus always integer multiples of = s. For our observations, the spectra were recorded every s. For calibration, the system periodically injects noise of K every (with s of noise on and s of noise off).
G165 VHVC was covered by eleven OTF fields, each with an angular size of 1∘ 1∘. The observation of each field is required to be completed within two hours to ensure a proper range of . Thus, we adopt a maximum scanning speed of 30′′ s-1 and a scanning row separation () of 2.7′ (Supplementary Fig. 1). To achieve the criteria of Nyquist sampling, the feed was rotated by 53.5∘ clockwise. The overhead time cost between two consecutive scanning rows is fixed to be 54 s. Each field requires an observing time of hours. In total, 13 hours were required to cover all the eleven fields.
At present, there is no widely used data reduction pipeline for the FAST OTF observations. Below follows the description of our data reduction procedure.
Gridding: At present, FAST only records the start time () of the OTF observations. The sampling time () of the spectrum can then be derived as . The accurate value (0.50331648 s) of must be adopted. If an approximate value of 0.5 s was adopted, the maximum deviation of the derived would be 26 s, leading to a maximum angular shift of . We then grid the data into pixels according to the targeting position derived from . The distribution of the effective integration time () of the final cube is shown in Extended Data Fig. 1.
Doppler correction: For each spectrum, Doppler corrections for the rotation of Earth and the movement of Earth and Sun were applied according to its coordinate and , using the public Python package PyAstronomy.
Calibration: The spectra are dominated by very strong radio frequency interference (RFI) within the frequency range 1.15–1.3 GHz, but are not significantly affected by RFI around 1.4 GHz. The spectra around 1.4 GHz were extracted for further reduction and analysis. For each data block of one minute observations, the average spectra of the data with the injected noise on and off were obtained and denoted as and , respectively. The spectral baselines () of and were obtained through polynomial fitting on line-free channels, and the difference between their spectral baselines is denoted as
Each spectrum (with the one denoted as ) in the data block was then calibrated through
Split HI cube: First-order polynomial fitting has been adopted to subtract the spectral baseline of the Hi cube. The spectral baseline is mainly composed of the system temperature () and the continuum emission (). We applied 2-d polynomial fitting on the point-source-free regions of the spectral baseline map to further separate and . The resultant is typically 19 K.
Robust check: The continuum image is dominated by background quasars (Supplementary Fig. 2). We applied 2- Gaussian fitting to these point-like sources with peak intensities larger than 3 K, yielding an FWHM of typically 3–3.4 arcmin, comparable to the beam size. The locations and fluxes of the point-like sources obtained by FAST (this work) and the NRAO VLA Sky Survey (NVSS)58 are consistent with each other (Supplementary Fig. 2), with a flux deviation smaller than several percent. We smoothed the Hi cube of FAST (this work) and of HI4PI 36 to the same resolution, and compared the spectra at a randomly chosen location (Supplementary 3). The two spectra are nearly identical. We conclude that the Hi cube we obtained is well calibrated and gridded.
2. Gaussian decomposition and filament extraction
The integrated intensity (moment 0) maps and channel maps of the FAST Hi cube reveal clear filamentary structures for all three (VHVC, HVC, and foreground) components of G165 (Fig. 1 and Extended Data Fig. 2). We apply Gaussian decomposition of the Hi spectra of G165 VHVC using scousepy and develop the tool ClimbRidge to extract the filamentary structures directly in the ppv cube.
Gaussian decomposition: We decomposed the velocity subcomponents of the G165 VHVC using scousepy, a widely used multi-component spectral line decomposition tool59, 60. At each pixel, G165 VHVC can be at most divided into three velocity subcomponents (Fig. 1). This procedure decomposed the ppv cube into multiple layers of the fitted (RA, DEC, ) points (Fig. 2).
Algorithm and code of ClimbRidge: ClimbRidge is a Python tool we have written to extract the filaments in -dimensional (-d) space following a gradient-ascent method – the density ridge algorithm50, 51. This algorithm models filaments in n- space as density ridges, where the eigenvector of the maximum eigenvalue of the Hessian matrix is parallel to the direction of the gradient. In practice, to find out the density ridges, a set of points are scattered in the -d space, and each point climbs along the gradient in the subspace of the other n-1 eigenvectors (those of all the eigenvalues except the maximum one). We provide both the CPU and GPU versions of ClimbRidge.
Test of ClimbRidge in 2-d case: In 2-d cases, the data at each pixel of the integrated intensity map are taken as the input of ClimbRidge to construct the density ridges. For comparison, we also apply FilFinder61 to find filaments on the 2-d map (Extended Fig. 3). The filaments found by the two methods are similar, and ClimbRidge is more sensitive to weak structures. FilFinder is a 2-d tool, while ClimbRidge can be directly applied to data of any dimension.
Extract filaments in 3-d ppv cube: In 3-d cases, the fitting results of scousepy are taken as the input of ClimbRidge. We reconstructed a smoothed ppv cube through convolving the fitting results of scousepy with a Gaussian kernel. The FWHM of the Gaussian kernal is 3.5′, 3.5′ and 15 km s-1 in the RA, DEC and velocity direction, respectively. In total, 15 000 points are scattered in the ppv space and move along the projected gradient. The resultant density ridges are shown in Fig. 1 and Extended Data Fig. 4.
MST analysis: We have developed a tool, fil3d, to analyse the ridge points produced by ClimbRidge, including (1) trimming out the isolated points, (2) minimal spanning tree analysis, (3) and find out the shortest path between two given points. We have built the minimal spanning trees (MSTs) to connect the density-ridge points (Extended Data Fig. 6), using the public Python package MiSTree, for both the 2-d and 3-d cases. To build the 3-d tree, the Euclidean distance is normalized by the units of (pixel, pixel, km s-1) in the three directions. A segment is defined as the path that connects two neighbouring nodes. A branch is defined as one or more connected segments. We also redistributed the density-ridge points through randomly rearranging their RA, DEC and , and then built a random tree. The maximuum segment length of the random tree is obviously shorter than that of the 3-d tree (Fig. 3). The 2-d and 3-d trees of G165 VHVC are qualitatively similar to each other (Extended Data Fig. 6). The lower end of the segment-length PDF of the 3-d tree is more flattened than that of the 2-d tree. It implies that the density ridges in the ppv cube can trace more short filamentary branches, e.g., in the crowded region of branch B3 (Extended Data Fig. 4).
3. Derive physical parameters Adopting a thickness of the G165 VHVC of 1.5∘ (the value comparable with its angular size) and a layer number of 4, the thickness of one layer () is estimated to be
The Hi emission of the G165 VHVC is weak and optically thin (Fig. 1). The Hi column density can be directly calculated through 62
The median value of the Hi integrated intensity is 37 K km s-1, yielding a median column density of cm-2. It corresponds to a typical global volume density, , of 0.2 cm-2. The linear mass () of the Hi filament can then be derived through
where is the distance between G165 VHVC and us. The gravity-related critical density of a filament can be estimated through 7
where is the gravitational constant. Inserting the typical values of Table 1 into Eq. 5, we get pc-1. Taking as 10 km s-1, we get pc-1, much larger than the value of . The total Hi flux of the G165 VHVC is 350 K km s -1 deg2, corresponding to a total mass of
where is the mass of hydrogen atom and is the angular coverage. The virial mass of the G165 VHVC is estimated to be
adopting and kpc. is also much larger than , implying that the self-gravity of G165 VHVC can be ignored.
4. Ideal MHD simulation
We use the multi-physics code ORION2 that is based on the Godunov scheme 63, 64 to simulate the observed G165 VHVC. The ideal MHD module in ORION2 is based on the dimensionally unsplit corner transport upwind scheme65 incorporating the constrained transport framework66, 67. ORION2 also has the adaptive refinement capability but we shall only use a single grid in our simulations here. From observation, the G165 VHVC is likely supersonic with a of 3. Besides this value, we have also performed a few more models at of 0.5 (subsonic), 1.0 (transonic), and 4.0 (even more supersonic) for comparison. We do not have magnetic field information of the region but we expect a magnetic field will affect the ionized WNM. Therefore, we perform a parameter study of with values 0.5, 1, and 2 using turbulence boxes with periodic boundary conditions. Here, with the Alfvén speed km s-1 for an Hi system. We assume that the VHVC is exposed to a uniform ultraviolet field and has a constant temperature. The overall volume density of G165 VHVC is smaller than the typical phase-transition density of WNM 43. The enhanced cooling rate in the most central regions of the filamentary hubs may lead to the existence of a colder HI phase, but it should not significantly influence the global dynamics. Therefore, we use the isothermal equation of state in the simulation. As discussed above, we also ignore gravity in the simulations. The simulated models are thus scaleless and only controlled by and . In each model, we start with an initially uniform density and magnetic field along the z-axis. We use the recipe of MacLow68 in driving the turbulence boxes with a pure solenoidal velocity field continuously at the largest scale (k=1-2) for two crossing times , where is the length of the box and is the root-mean-square turbulence velocity. For , is the sound speed . For comparison with observation, we only use the simulation data at .
To mimic a random observational direction, we inspected the simulated models at the viewing angle of from the initial magnetic field orientation (z-axis), and produced the simulated ppv cubes under the optically thin limit (Supplementary Fig. 4). Further, we smoothed the simulated ppv cubes to have a map size (256 pixel) similar to that of observation. Gaussian decomposition was applied to the simulated ppv cubes, as to the observed data. For the best-fit model ( and ), we further extracted its filamentary structures in the ppv space using ClimbRidge (Extended Data Fig. 7).
Data Availability Owing to data size, the reduced data used for this study are available from the corresponding authors upon reasonable request.
Code Availability
The FAST OTF data reduction script we developed for this work can found at https://gitee.com/liuxunchuan/fast_otf_datareduct. The filament extracting code we developed for this work, ClimbRidge and fil3d, can be found at https://gitee.com/liuxunchuan/sometools/tree/master/filament/.
The MST analysis tool is available at https://pypi.org/project/mistree/. The compared filament extracting method FilFinder can be found at https://github.com/e-koch/FilFinder. Image processing package scikit-image can be found at https://pypi.org/project/scikit-image/. The ORION2 code is available at https://joss.theoj.org/papers/10.21105/joss.03771.
Acknowledgements We wish to thank the staff of the FAST for their kind help provided during the observations. This work has been supported by the National Key R&D Program of China (No. 2022YFA1603100) and National Natural Science Foundation of China (NSFC) through grants (No. 12203086, No.12073061, No.12122307 and No. 12033005). X.L. acknowledges the supports by CPSF No. 2022M723278. T.L. and P.S.Li acknowledge the supports by the Tianchi Talent Program of Xinjiang Uygur Autonomous Region and the international partnership program of Chinese Academy of Sciences through grant No.114231KYSB20200009. This research was carried out in part at the Jet Propulsion Laboratory, which is operated by the California Institute of Technology under a contract with the National Aeronautics and Space Administration (80NM0018D0004). We thank the anonymous reviewers for providing feedback that helps to solidify this work.
Author contributions
X.C. L. was the principal investigator of the FAST project. He led the data reduction, filament-extracting code development, scientific analysis and wrote the manuscript. T. L. contributed to the FAST proposal, the interpretation of results and writing of the manuscript. P.S. L. conducted the MHD simulation. All authors participated in the discussion of results, and revision of the manuscript.
Competing interests
The authors declare no competing interests.
Filament | length | V | |||
(arcmin) | (arcmin) | (km s-1) | (km s-1 pc-1) | (K) | |
B1 | 90 | 6.1 | 25.3 | -0.02 | 0.73 |
B2 | 54 | 9.5 | 23.9 | 0.1 | 0.85 |
B3 | 60 | 7.0 | 29.4 | -0.1 | 0.40 |
Note: The listed parameters include the angular length, velocity gradient (), angular width (), velocity width (), and the averaged brightness temperature (). Adopting a distance of 15 kpc, the converting factor between the linear scale (pc) and angular scale (arcmin) is 4.36 pc arcmin-1.